The length of a rectangle is 6 inches more than width. The perimeter of the rectangle can be no more than 48 inches. What is maximum width?
Answer by ikleyn(52786)

.

    x + (x+6) <= 48/2

      2x + 6  <= 24

      2x      <= 24 - 6

      2x      <=   18

       x      <=   18/2 = 9.


ANSWER.  The maximum width is 9 inches.

Solved.